2. [6 pts.] The gross domestic product (GDP) of a developing country from 1999 (t=0) to
2007 is approximated by
f(t) = −0.2t^3 + 2.4t^2 + 60 (0 <_ t _> 8)

where f(t) is measured in billions of dollars. Find the maximum growth rate of the
country’s GDP in the given period.
Thank you so much for your time. I have been trying to do this problem for a couple of days now. I work full time and go to school full time so its hard to keep up in class.

Answer by stanbon(75887)

The gross domestic product (GDP) of a developing country from 1999 (t=0) to
2007 is approximated by f(t) = −0.2t^3 + 2.4t^2 + 60 (0 <_ t =< 8)
where f(t) is measured in billions of dollars. Find the maximum growth rate of the country’s GDP in the given period.
-------------------------------------
You could do this in two ways:
Using a graphing calculator:
When t=8, f(8) = $111.2 billion
-------------------------------------
If you are studying calculus, take the derivative and set to zero.
-0.6t^2+4.8t = 0
-0.6t(t-8) = 0
t = 8,
Then f(8) = $111.2 billion
================================
Cheers,
Stan H.